OpenStudy (anonymous):

Show that \[||x||_{\infty}\leq||x||_2\leq ||x||_1 \forall x\in \mathbb R^n\]

OpenStudy (anonymous):

induction?

OpenStudy (jamesj):

No ... instead consider the function in p \[f(x; p) = \left( \sum_{i=1}^n |x_i|^p \right)^{1/p}\] Now for a given x, show first that: \[||x||_{\infty} \leq ||x||_p = f(x; p)\] In fact it is also possible to show--but not necessary here to do so--that \[\lim_{p \rightarrow \infty} ||x||_p = ||x||_{\infty}\]

OpenStudy (anonymous):

this is what i want to show. how does defining the function help?

OpenStudy (jamesj):

Because now you just need to differentiate f with respect to p and show it has the right sign.

OpenStudy (jamesj):

partially of course.

OpenStudy (anonymous):

ii will try it. thank you again

OpenStudy (jamesj):

Try it; it's slightly messy as you need to consider a few different cases but you'll get the general result. However, for your problem, it might be easier to consider this problem directly. E.g., ||x||_inf = max | xi | = |xj | say. Then ||x||_inf = | xj | \leq ||x||2 iff |xj|^2 \leq |x1|^2 + |x2|^2 + .... + |xj|^2 + ... + |xn|^2, which is obvious That's one part of the inequality. Try the other inequality directly and I think you won't have a hard time convince yourself that ||x||2 \leq ||x||1